OpenStudy (anonymous):

Solve the following equation: 7 + x + 2 = 9 + 3

OpenStudy (anonymous):

i think it is 3

OpenStudy (anonymous):

X = 3

OpenStudy (gorv):

9+x=9+3 x=3

OpenStudy (anonymous):

thanks guys :)